• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:03:41 on October 25, 1998

    Entry number 3590

    keyword=Beam Time Accounting
    Hall A

    10/25/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: Deur A.
    Shift Workers: Brown D., Kramer K., Zhang B.

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    1 1 1 1 40 1 1 55 7.58333333333333
    Beam Available
    but Not in Use
    0 0 0 0 20 0 0 0 0.333333333333333
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 5 0.0833333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 40 1 1 1 7.66666666666667
    Planned Config-
    uration Change
    0 0 0 0 20 0 0 0 0.333333333333333
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 20 0 0 0 0.333333333333333
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: E=3.4 GeV, P=3.028 GeV/c, target in transverse configuration.production runs 8:50 Asked MCC for Bscope+checked raster size. Good 8:56 target Ion chamber readout: 106.2 Rad (total dose) 12:00 NMR measurement (have to switch from |_ to \\ , then NMR, then back to |_. Target polarization 28 % (loss). Used new calibration from Gordon 14:07 coda crashed (100 % deadtime, o Hz)]

    Data Entry by: Deur A.